# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v1

@@
Abstract
Decision making is a central problem in AI that can be formalized using a Markov Decision Process. A problem is that, with increasing numbers of (indistinguishable) objects, the state space grows exponentially. To compute policies, the state space has to be enumerated. Even more possibilities have to be enumerated if the size of the action space depends on the size of the state space, especially if we allow concurrent actions. To tackle the exponential blow-up in the action and state space, we present a first-order representation to store the spaces in polynomial instead of exponential size in the number of objects and introduce Foreplan, a relational forward planner, which uses this representation to efficiently compute policies for numerous indistinguishable objects and actions. Additionally, we introduce an even faster approximate version of Foreplan. Moreover, Foreplan identifies how many objects an agent should act on to achieve a certain task given restrictions. Further, we provide a theoretical analysis and an empirical evaluation of Foreplan, demonstrating a speedup of at least four orders of magnitude.
@@ Abstract
8114
1 Introduction
Decision making problems can be formalized using Markov Decision Processes (MDPs). To compute a policy for an MDP, the state and action spaces have to be enumerated to find the best possible action for each state. But, with increasing number of (indistinguishable) objects, the state space grows exponentially. In case the actions depend on the states or objects as well, even more possibilities have to be enumerated, let alone if further concurrent actions are allowed. For concurrent actions, all possible action combinations have to be enumerated, yielding an exponentially-sized action space, which is why concurrency is rarely modelled. However, consider the following example: A small town is haunted by an epidemic. To fight the epidemic, the town’s mayor can impose travel bans on the town’s citizens. Certainly, the mayor can confine all citizens to their homes, stopping the epidemic. However, the citizens’ overall welfare is important as well. Therefore, the mayor is interested in the best decision of imposing travel bans (concurrent actions) w.r.t. confining the epidemic, while keeping the citizens’ welfare above a certain threshold. In addition, there can be circumstances, where the policy has to be adapted, e.g., if a wave of infections is on the rise, threatening to infect the majority of the population. Computing this problem on a propositional level, with every citizen explicitly represented, like approaches such as MDPs do, blows up the state and action space, as it requires exponential enumeration of all subsets of the population. However, there are groups of citizens behaving identically w.r.t. getting sick (and well again) as well as regarding their welfare if a travel ban is imposed, i.e., these citizens are indistinguishable for the mayor. Within these groups, it does not matter on which exact citizen the mayor imposes a travel ban on, only on how many she imposes a travel ban on. Additionally, many computations over subsets of indistinguishable citizens are redundant. Thus, we propose to drastically reduce the search and action space by using a first-order representation, grouping indistinguishable citizens, and use this representation to plan group-wise by reasoning about a representative and then projecting the result to the whole group.
Contribution
First, to be able to represent numerous indistinguishable objects and actions, we use probabilistic relational models as a representation in (factored) MDPs (fMDPs), which yields relational factored MDPs (rfMDPs). Second, we propose Foreplan to carry out planning in rfMDPs. State of the art algorithms run in time exponential in the number of objects. To reduce the time taken to polynomial in the number of objects, we define the relational cost graph, which is used by Foreplan to encode the state and action space in an efficient way. Denoting the number of cliques by $c$ and the size of the largest clique by $w$ in the relational cost graph, Foreplan runs exponential in $c$ and $w$. Both parameters are structural ones, mostly small and fixed, and thus independent of the number of objects. Therefore, Foreplan is an efficient planner w.r.t the number of objects. Third, using approximation, we can reduce the runtime even further. We propose Approximate Foreplan, whose runtime is polynomial in $c$. For Approximate Foreplan, we show a speedup of at least four orders of magnitude. Fourth, we show that Foreplan efficiently identifies how many objects the agent should act on to achieve a task given certain restrictions. The resulting extension of Foreplan runs in polynomial time in the number of objects, when the underlying relational cost graph has bounded maximal clique size and if the model is liftable and the evidence for state and action is liftable as well.
Related Work
Bellman [2] introduces MDPs, which Boutilier et al. [4] extend to fMDPs by factorizing the transition function. Factorizing also the value function, Guestrin et al. [16] provide two approximate algorithms for solving planning in fMDPs. Dean et al. [10] cluster the state space of fMDPs to reduce the state space even further. Givan et al. [15] group equivalent states based on the notion of bisimulation. Both approaches lack the ability to handle concurrent actions efficiently. MDPs can be generalized to partially observable MDPs, in which the agent is uncertain in which state the environment currently is [19]. Sanner and Kersting [29] add the first-order perspective to partially observable MDPs, focusing on the observations and without concurrent actions. Bernstein et al. [3] extend partially observable MDPs to have a set of decentralized agents. Braun et al. [6] work with groups of indistinguishable agents. This is similar to our approach, in which we handle sets of indistinguishable state variables. However, they do not provide a solution environment.
The idea of lifting is to carry out computations over representatives for groups of indistinguishable random variables [26, 20, 9]. There are online decision making approaches adding action and utilitity nodes to this representation [1, 14, 13], here, we focus on offline planning. To carry out even more lifted computations, Taghipour [30] extends lifted probabilistic inference by a generalized counting framework, which we extend later on. Using a first-order representation for states, actions and objects, Boutilier et al. [5] exploit the relational structure for planning using MDPs, still without concurrent actions. To specify factored transition models in first-order MDPs, Sanner and Boutilier [28] introduce factored first-order MDPs, which uses a backward search. While they consider actions on subsets of objects, they still treat every object individually. We apply lifting for efficient handling of the objects. Moreover, we prove on which models our algorithm runs efficiently. For a survey on planning using first-order representations, we refer to Corrêa and De Giacomo [7].
Structure
The remainder of this paper is structured as follows: First, we present preliminaries for (factored) MDPs, including lifting. Then, we introduce rfMDPs to group indistinguishable objects. Afterwards, we propose Foreplan, which exploits these indistinguishable objects using a compact state representation, to efficiently support decision making with concurrent actions. We also show how Foreplan identifies how many objects to act on for achieving a certain task. Further, we provide a theoretical analysis and empirical evaluation of Foreplan.
2 Preliminaries
2.1 (Factored) Markov Decision Processes
Definition 1 (Markov Decision Process).
@@ 2 Preliminaries
A Markov Decision Process is a tuple $(\mathbf{S},\mathbf{A},T,R)$ with a set of states $\mathbf{S}$, a set of actions $\mathbf{A}$, a reward function $R:\mathbf{S}\mapsto\mathbb{R}$ and a transition function $T:\mathbf{S}\times\mathbf{A}\times\mathbf{S}\mapsto[0,1]$. The value $T(s,a,s^{\prime})$ is the probability $P(s^{\prime}\mid s,a)$ of transitioning from state $s\in\mathbf{S}$ to state $s^{\prime}\in\mathbf{S}$ if action $a\in\mathbf{A}$ is performed.
@@
The rewards are additive, possibly discounted by a factor $\gamma\in[0,1]$. An MDP is fully observable and has the first order Markov property, i.e., the probability of the next state only depends on the current state and action. Let us have a look at a simple, (incomplete) example of an MDP.
Example 1.
@@
Suppose we have the states healthy and sick. Our agent has two possible actions: Travelling or staying at home. When the agent is in state sick and travels, she stays sick with a probability of $0.9$. The agent obtains a reward of $-1$ if sick and $1$ if healthy.
@@
To find the utility of a state algorithmically, we find a value function $V$ satisfying the Bellman equation. The value function induces a policy by selecting the action that yields the maximum expected value. For computing a value function, we can use a linear programming formulation [12, 27]:
$\begin{array}[]{ll@{}}\text{Variables:}&V(s)\quad\forall s\in\mathbf{S}\ ;\\ \text{Minimize:}&\displaystyle\sum_{s\in\mathbf{S}}\alpha(s)V(s)\ ;\\ \text{Subject to:}&\forall s\in\mathbf{S},\ \forall a\in\mathbf{A}:\\ &V(s)\geq R(s)+\gamma\displaystyle\sum_{s^{\prime}\in\mathbf{S}}P(s^{\prime}\mid s,a)\cdot V(s^{\prime}),\\ \end{array}$ (2)
where the coefficients $\alpha(s)$ are arbitrary positive numbers, e.g., an equal distribution over all states [27]. Planning in MDPs can be solved in polynomial time w.r.t to the state space size [25]. But what if the state space becomes very large, e.g., exponential in the number of objects? For retaining an efficient transition model, fMDPs make use of state variables for the objects. The state space is then spanned by the state variables. In our paper, all state variables are Boolean, but can be easily extended to non-Boolean.
Definition 3 (Factored MDP).
A factored MDP is a tuple $(\mathbf{S}^{\prime},\mathbf{A},T,R)$, where $\mathbf{S}^{\prime}$ contains $m$ state variables $S_{1},\dots,S_{m}$ with Boolean cardinality. Accordingly, the state space is $\mathbf{S}=\{0,1\}^{m}$. The transition function $T$ is, for each action, factored according to a (Dynamic) Bayesian network:
@@
where $Pa(S_{i}^{\prime})\subseteq\mathbf{S}$ denotes the set of parents of $S_{i}^{\prime}$ in the Bayesian network, $S$ the old state, $S^{\prime}$ the new state and $S_{i}^{\prime}$ the $i$-th state variables in the respective state. For a given state $s\in\mathbf{S}$, we denote with $s_{i}$ the assignment of state variable $S_{i}$ in state $s$.
For handling numerous objects, we use parameterized graphical models, which we introduce in the next subsection.
2.2 Parameterized Graphical Models
Often, we have indistinguishable random variables, leading to the same computations. We can tackle redundant computations by parameterizing our probabilistic model and grouping indistinguishable variables, so that inference in the probabilistic model becomes tractable w.r.t. domain sizes by using representatives during calculations [24]:
Definition 4 (Parfactor model [31]).
Let $\mathbf{W}$ be a set of random variable names, $\mathbf{L}$ a set of logical variable (logvar) names, $\Phi$ a set of factor names, and $\mathbf{D}$ a set of constants (universe). All sets are finite. Each logvar $L$ has a domain $\mathcal{D}(L)\subseteq\mathbf{D}$. A constraint $C$ is a tuple $(\mathcal{X},C_{\mathcal{X}})$ of a sequence of logvars $\mathcal{X}=(X_{1},\dots,X_{n})$ and a set $C_{\mathcal{X}}\subseteq\times_{i=1}^{n}\mathcal{D}(X_{i})$. The symbol $\top$ for $C$ marks that no restrictions apply, i.e., $C_{\mathcal{X}}=\times_{i=1}^{n}\mathcal{D}(X_{i})$. A parameterized random variable (PRV) $B(L_{1},\dots,L_{n}),n\geq 0,$ is a syntactical construct of a random variable $B\in\mathbf{W}$ possibly combined with logvars $L_{1},\dots,L_{n}\in\mathbf{L}$. If $n=0$, the PRV is parameterless and constitutes a propositional random variable (RV). The term $\mathcal{R}(B)$ denotes the possible values (range) of a PRV $B$. An event $B=b$ denotes the occurrence of PRV $B$ with range value $b\in\mathcal{R}(B)$. We denote a parameterized factor (parfactor) $g$ by $\phi(\mathcal{B})_{|C}$ with $\mathcal{B}=(B_{1},\dots,B_{n})$ a sequence of PRVs, $\phi:\times_{i=1}^{n}\mathcal{R}(B_{i})\mapsto\mathbb{R}^{+}$ a potential function with name $\phi\in\Phi$, and $C$ a constraint on the logvars of $\mathcal{B}$. A set of parfactors forms a model $G:=\{g_{i}\}_{i=1}^{n}$. With $Z$ as normalizing constant, $G$ represents the full joint distribution $P_{G}=\frac{1}{Z}\prod_{f\in gr(G)}f$, with $gr(G)$ referring to the groundings of $G$ w.r.t. given constraints. A grounding is the instantiation of each parfactor with an allowed constant.
Let us illustrate the definition of a parfactor model:
Example 2.
Let $\mathbf{W}=\{Sick,Epidemic\}$, $\mathbf{L}=\{M\}$, $\mathcal{D}(M)=\mathbf{D}=\{a,b,c,d,e,f,g,h\}$ with Boolean-valued PRVs $Sick(M)$ and $Epidemic$. Let $g(Sick(M),Epidemic)$ be the parfactor with potential $\phi$ defining the probability to be sick for all persons $\mathbf{D}$, given there is an epidemic (or not). The grounded model then consists of the eight factors $\phi(Sick(a),Epidemic)$, $\phi(Sick(b),Epidemic)$, …, $\phi(Sick(h),Epidemic)$.
Next, we present rfMDPs, integrating a parfactor model.
3 Relational Factored MDPs
Now, we present a first-order representation for indistinguishable objects and numerous actions on collections of them. Let us generalize Example 1 to an arbitrary number of persons behaving in the same way as the agent in Example 1:
Example 3 (Epidemic).
There is a set $\mathbf{D}$ of persons living in a small town, represented by the logvar $M$ with $\mathcal{D}(M)=\mathbf{D}$. Each person can be sick or healthy, leading to the PRV $Sick(M)$. The government gets a bonus of $1$ for each healthy person and a penalty of $-1$ for each sick person. To combat an epidemic, the government can impose travel bans on persons, resulting in the action $Restrict(M)$ to impose a travel ban on a subset of persons. Moreover, each person can travel or not, leading to the PRV $Travel(M)$. The government gets a bonus of $2$ for each person travelling. The PRV $Epidemic$ is influenced by the number of people travelling and influences the sickness of each person. Figure 1 shows the transition model for this example.
@@ 3 Relational Factored MDPs
Figure 1: Lifted representation of the transition model for Example 3. We abbreviate by using only the first letter(s) for each symbol.
Since an action can be applied to each person concurrently, the amount of possible actions is exponential due to the power set, i.e., all possible combinations for all persons. To prevent the exponential explosion, we introduce rfMDPs to group objects and actions. For modeling indistinguishable objects, we require all functions used in the rfMDP to be oblivious about the exact object, i.e., permuting the objects does not alter the output, and ensure that the functions work on representatives for each group. We start by defining rfMDPs as a mixture of fMDPs and a parfactor model. Afterwards, we investigate and define the effects on the actions and rewards.

### v2

@@
Abstract.
When allowing concurrent actions in Markov Decision Processes, whose state and action spaces grow exponentially in the number of objects, computing a policy becomes highly inefficient, as it requires enumerating the joint of the two spaces. For the case of indistinguishable objects, we present a first-order representation to tackle the exponential blow-up in the action and state spaces. We propose Foreplan, an efficient relational forward planner, which uses the first-order representation allowing to compute policies in space and time polynomially in the number of objects. Thus, Foreplan significantly increases the number of planning problems solvable, in an exact manner, in reasonable time, which we underscore with a theoretical analysis. To speed up the computations even further, we also introduce an approximate version of Foreplan, including guarantees on the error. Further, we provide an empirical evaluation of both Foreplan versions, demonstrating a speedup of several orders of magnitude. For the approximate version of Foreplan, we also empirically show that the induced error is often negligible.
Key words and phrases:
Planning, Markov Decision Process, Lifting, Probabilistic Graphical Models
1. Introduction
To compute a policy for a Markov decision processes (MDPs), combinations of the state and action spaces have to be enumerated to find the best possible action for each state. With increasing number of (indistinguishable) objects, the state space grows exponentially, leading to drastically more combinations to be enumerated. In case the actions also depend on the states or objects, the action space grows as well, leading to even more combinations for enumeration. Further, if one allows concurrent actions, all possible combinations of the actions, yielding an exponentially-sized action space, of the already enlarged action space have to be used during the enumeration. Therefore, given the number of possibilities that have to be enumerated, concurrency is rarely modelled. However, consider the following example: A small town is haunted by an epidemic. To fight the epidemic, the town’s mayor can impose travel bans on the town’s citizens. Certainly, the mayor can confine all citizens to their homes, stopping the epidemic. However, the citizens’ overall welfare is important as well. Therefore, the mayor is interested in the best decision of imposing travel bans (concurrent actions) w.r.t. confining the epidemic, while keeping the citizens’ welfare above a certain threshold. In addition, there can be circumstances where the policy has to be adapted, e.g., if a wave of infections is on the rise, threatening to infect the majority of the population. Computing this problem on a propositional level, with every citizen explicitly represented, like approaches such as MDPs do, blows up the state and action space, as it requires exponential enumeration of all subsets of the population. However, there are groups of citizens behaving identically w.r.t. getting sick (and well again) as well as regarding their welfare if a travel ban is imposed, i.e., these citizens are indistinguishable for the mayor. Within these groups, it does not matter on which exact citizen the mayor imposes a travel ban on, only on how many she imposes a travel ban on. Additionally, many computations over subsets of indistinguishable citizens are redundant. Thus, we propose to drastically reduce the search and action space by using a first-order representation, grouping indistinguishable citizens, and use this representation to plan group-wise by reasoning about a representative and then projecting the result to the whole group.
Contribution
First, to be able to represent numerous indistinguishable objects and actions, we use probabilistic relational models as a representation in (factored) MDPs (fMDPs), which yields relational factored MDPs (rfMDPs). Second, as our key contribution, we propose Foreplan to carry out exact planning in rfMDPs. Foreplan uses our first-order representation to keep the action and state space only polynomial in the number of objects, which we prove by using our so-called relational cost graph. Foreplan remains exponential in the number of cliques $c$ and size $w$ of the largest clique in the relational cost graph, but both parameters are mostly small and independent of the domain sizes. Therefore, Foreplan is an efficient planner w.r.t the number of objects. We show a speedup of at least four orders of magnitude is achievable. Last, following the approximation ideas of approx_factored_mdps, we can reduce the runtime even further. We propose Approximate Foreplan, whose runtime is polynomial in $c$. For Approximate Foreplan, we show a speedup of at least four orders of magnitude.
Related Work
mdp introduces MDPs, which factored_mdp extend to fMDPs by factorizing the transition function. Factorizing also the value function, approx_factored_mdps provide two approximate algorithms for solving planning in fMDPs. dean2013model cluster the state space of fMDPs to reduce the state space even further. givan2003equivalence group equivalent states based on the notion of bisimulation. Both approaches lack the ability to handle concurrent actions efficiently. MDPs can be generalized to partially observable MDPs, in which the agent is uncertain in which state the environment currently is kaelbling1998planning. sanner2010symbolic add the first-order perspective to partially observable MDPs, focusing on the observations and without concurrent actions. bernstein2002complexity extend partially observable MDPs to have a set of decentralized agents. braun2022lifting work with groups of indistinguishable agents. This is similar to our approach, in which we handle sets of indistinguishable state variables. However, they do not provide an algorithm.
The idea of lifting is to carry out computations over representatives for groups of indistinguishable random variables lve; kersting2012lifted; raedt2016statistical. There are online decision making approaches adding action and utilitity nodes to this representation apsel2011extended; 10.1007/978-3-030-12738-1_10; 10.1007/978-3-030-18305-9_33, here, we focus on offline planning. To carry out even more lifted computations, crv extends lifted probabilistic inference by a generalized counting framework, which we extend later on. Using a first-order representation for states, actions and objects, boutilier2001symbolic exploit the relational structure for planning using MDPs, still without concurrent actions. To specify factored transition models in first-order MDPs, sanner2007approximate introduce factored first-order MDPs, which uses a backward search and provide only an approximate algorithm. In contrast, we propose an exact algorithm by applying lifting. Later, we also introduce an approximate version. Moreover, we prove on which models our algorithm runs efficiently. For a survey on planning using first-order representations, we refer to correa2024lifted.
Structure
The remainder of this paper is structured as follows: First, we present preliminaries for (factored) MDPs, including lifting. Then, we introduce rfMDPs to group indistinguishable objects. Afterwards, we propose Foreplan, which exploits these indistinguishable objects using a compact state representation, to efficiently support decision making with concurrent actions. Further, we provide a theoretical analysis and empirical evaluation of Foreplan.
2. Preliminaries
2.1. (Factored) Markov Decision Processes
@@ 2. Preliminaries
A Markov Decision Process is a tuple $(\mathbf{S},\mathbf{A},T,R)$ with a set of states $\mathbf{S}$, a set of actions $\mathbf{A}$, a reward function $R:\mathbf{S}\mapsto\mathbb{R}$ and a transition function $T:\mathbf{S}\times\mathbf{A}\times\mathbf{S}\to[0,1]$. The value $T(s,a,s^{\prime})$ is the probability $P(s^{\prime}\mid s,a)$ of transitioning from state $s\in\mathbf{S}$ to state $s^{\prime}\in\mathbf{S}$ if action $a\in\mathbf{A}$ is performed.
@@
The rewards are additive, possibly discounted by a factor $\gamma\in[0,1)$. An MDP is fully observable and has the first order Markov property, i.e., the probability of the next state only depends on the current state and action. Let us have a look at a simple, (incomplete) example of an MDP.
[…]
@@
Suppose we have the states healthy and sick. Our agent has two possible actions: Travelling or staying at home. When the agent is in state sick and travels, she stays sick with a probability of $0.9$. The agent obtains a reward of 1-1 if sick and $1$ if healthy.
@@
To find the utility of a state algorithmically, we find a value function $V$ satisfying the Bellman equation. The value function induces a policy by selecting the action that yields the maximum expected value. For computing a value function, we can use a linear programming formulation mdp_lp_orig; mdp_lp:
[…]
where the coefficients $\alpha(s)$ are arbitrary positive numbers, e.g., an equal distribution over all states mdp_lp. Planning in MDPs can be solved in polynomial time w.r.t to the state space size papadimitriou1987complexity. But what if the state space becomes very large, e.g., exponential in the number of objects? For retaining an efficient transition model, fMDPs make use of state variables for the objects. The state space is then spanned by the state variables and is exponential in the number of state variables. For simplicity, all state variables are Boolean, but can be easily extended to non-Boolean.
[…]
A factored MDP is a tuple $(\mathbf{S}$, $\mathbf{A}$, $T$, $R)$, where the state space is the cross product $\mathbf{S}=\{0,1\}^{m}$ of the sub-spaces corresponding´to the state variables $S_{1},\dots,S_{m}$ with Boolean range. The transition function $T$ is, for each action, factored:
@@
where $Pa(S_{i}^{\prime})\subseteq\mathbf{S}$ denotes the set of parents of $S_{i}^{\prime}$ in the previous state, $S$ the old state, $S^{\prime}$ the new state and $S_{i}^{\prime}$ the $i$-th state variables in the respective state. For a given state $s\in\mathbf{S}$, we denote with $s_{i}$ the assignment of state variable $S_{i}$ in state $s$.
When having numerous objects, the state space explodes. But when these objects are indistinguishable, we can use parameterized graphical models to encode the state space and transition model compactly.
2.2. Parameterized Graphical Models
Having indistinguishable random variables leads to the redundant computations. We can tackle redundant computations by parameterizing our probabilistic model and grouping indistinguishable variables, so that inference in the probabilistic model becomes tractable w.r.t. domain sizes by using representatives during calculations niepert2014tractability. We first define parameterized random variables (PRVs) to group same behaving variables:
@@ 2.2. Parameterized Graphical Models
Let $\mathbf{W}$ be a set of random variable names, $\mathbf{L}$ a set of logical variable (logvar) names, and $\mathbf{D}$ a set of constants (universe). All sets are finite. Each logvar $L$ has a domain $\mathcal{D}(L)\subseteq\mathbf{D}$. A constraint $C$ is a tuple $(\mathcal{X},C_{\mathcal{X}})$ of a sequence of logvars $\mathcal{X}=(X_{1},\dots,X_{n})$ and a set $C_{\mathcal{X}}\subseteq\times_{i=1}^{n}\mathcal{D}(X_{i})$. The symbol $\top$ for $C$ marks that no restrictions apply, i.e., $C_{\mathcal{X}}=\times_{i=1}^{n}\mathcal{D}(X_{i})$. A PRV $B(L_{1},\dots,L_{n}),n\geq 0,$ is a syntactical construct of a random variable $B\in\mathbf{W}$ possibly combined with logvars $L_{1},\dots,L_{n}\in\mathbf{L}$. If $n=0$, the PRV is parameterless and constitutes a propositional random variable (RV). The term $\mathcal{R}(B)$ denotes the possible values (range) of a PRV $B$. An event $B=b$ denotes the occurrence of PRV $B$ with range value $b\in\mathcal{R}(B)$.
We create a PRV for the health status of each person in the town:
[…]
Let $\mathbf{W}=\{Sick,Epidemic\}$, $\mathbf{L}=\{M\}$, $\mathcal{D}(M)=\mathbf{D}=\{a,b,c,d,e,f,g,h\}$ with Boolean-valued PRVs $Sick(M)$ and $Epidemic$.
EpidemicEpidemic is a parameterless PRV. To set PRVs into relation, we use parameterized factors:
[…]
Let $\Phi$ be a set of factor names. We denote a parameterized factor (parfactor) $g$ by $\phi(\mathcal{B})_{|C}$ with $\mathcal{B}=(B_{1},\dots,B_{n})$ a sequence of PRVs, $\phi:\times_{i=1}^{n}\mathcal{R}(B_{i})\to\mathbb{R}^{+}$ a potential function with name $\phi\in\Phi$, and $C$ a constraint on the logvars of $\mathcal{B}$. A set of parfactors forms a model $G:=\{g_{i}\}_{i=1}^{n}$. With $Z$ as normalizing constant, $G$ represents the full joint distribution $P_{G}=\frac{1}{Z}\prod_{f\in gr(G)}f$, with $gr(G)$ referring to the groundings of $G$ w.r.t. given constraints. A grounding is the instantiation of each parfactor with an allowed constant.
Continuing Example 2.2, let us define the parfactor model:
[…]
Let $\forall m\in M:\phi(Sick(m),Epidemic)_{\top}$ be the parfactor with potential $\phi$ defining the probability to be sick for all persons $\mathbf{D}$, given there is an epidemic (or not). The grounded model then consists of the eight factors $\phi(Sick(a),Epidemic)$, $\phi(Sick(b)$, $Epidemic)$, …, $\phi(Sick(h),Epidemic)$ compared to one parfactor in the lifted model. Table 1 shows the potential function.
Table 1. Potential function giving the probability of a person being sick given whether there is an epidemic.
EpidemicEpidemic $Sick(M)$
0 0 0.9
0 1 0.1
1 0 0.3
1 1 0.7
Next, we integrate the parfactor into fMDPs to achieve a compact representation of the state and action spaces.
3. Relational Factored MDPs
We present the first-order representation for the state and action space as well as the transition function, which allows to efficiently represent indistinguishable objects as well as concurrent action, which then in turn can be exploited by an appropriate algorithm. To that end, we present Foreplan in the next section.